OpenStudy (anonymous):

if \left| x _{1} -x _{2}\right|=sqrt[+]{(x _{1}-x _{2})^2} Is the something similar for \left| x _{1} +x _{2}\right|

OpenStudy (anonymous):

\[\left| x _{1} -x _{2}\right|=\sqrt[+]{(x _{1}-x _{2})^2} \] ?similar to \[\left| x _{1} +x _{2}\right|\]

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

Either way, you're just taking absolute value of a number which is always positive, so it should be equal to the positive square-root of its square.